NPP Projection Close 1/27
To do this evening/ was completed:
Presentation Issues:
Preshow Clips/Playlist: Updated the "TLR_SE_PrecodeParade_FEB_DancingLady" clip to its newest version ("V2") in all necessary pre-shows this week and next per Kat's request
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Let the Right One In was delivered this evening and brought up to booth 1
Outgoing/Shipping (content): Storage space across the servers, especially in 2, 4, and 5, is extremely limited. There is not much to remove since there are so many first runs being shuffled around. There are currently no issues with playback but please keep an eye out in the week(s) ahead as titles come and go.
Presentation Issues:
Preshow Clips/Playlist: Updated the "TLR_SE_PrecodeParade_FEB_DancingLady" clip to its newest version ("V2") in all necessary pre-shows this week and next per Kat's request
Built CCAP and OCAP show playlists for The Moment and saved them to the LMS (note: the CCAP version has a 7.1 mix which I've added to its playlist but the OCAP is only in 5.1)
Trailer Updates:
Special Events/ AV set up: A Few Good Men was not tech'd/did not have a volume cue so I started the feature at 7.0 and adjusted accordingly. Otherwise the screening went well. All reels are tails out on the work bench in booth 1 and can be re-inspected tomorrow. Both projectors were cleaned and the exhausts shut after the show + Possessor: Uncut screening in theater 4 went smoothly
Trailer Updates:
Special Events/ AV set up: A Few Good Men was not tech'd/did not have a volume cue so I started the feature at 7.0 and adjusted accordingly. Otherwise the screening went well. All reels are tails out on the work bench in booth 1 and can be re-inspected tomorrow. Both projectors were cleaned and the exhausts shut after the show + Possessor: Uncut screening in theater 4 went smoothly
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Let the Right One In was delivered this evening and brought up to booth 1
Outgoing/Shipping (content): Storage space across the servers, especially in 2, 4, and 5, is extremely limited. There is not much to remove since there are so many first runs being shuffled around. There are currently no issues with playback but please keep an eye out in the week(s) ahead as titles come and go.
Removed Possessor: Uncut feature content from the LMS and theater 4 following the show
Print/video Inspection: Started inspecting Let the Right One In (reels 1-3) and uploaded some stills to the shared drive
Asset tracking doc updated/checked?:
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Yes
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?: Yes
Projection cart by entrance?: Yes
Additional Notes: Updated the first run dry erase board in 4/6 for the week of 1/30 (including re-releases)
Print/video Inspection: Started inspecting Let the Right One In (reels 1-3) and uploaded some stills to the shared drive
Asset tracking doc updated/checked?:
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Yes
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?: Yes
Projection cart by entrance?: Yes
Additional Notes: Updated the first run dry erase board in 4/6 for the week of 1/30 (including re-releases)
Comments
Post a Comment